# DINeuro: Distilling Knowledge from 2D Natural Images via Deformable Tubular Transferring Strategy for 3D Neuron Reconstruction

### v1

@@ Abstract
Reconstructing neuron morphology from 3D light microscope imaging data is critical to aid neuroscientists in analyzing brain networks and neuroanatomy. With the boost from deep learning techniques, a variety of learning-based segmentation models have been developed to enhance the signal-to-noise ratio of raw neuron images as a pre-processing step in the reconstruction workflow. However, most existing models directly encode the latent representative features of volumetric neuron data but neglect their intrinsic morphological knowledge. To address this limitation, we design a novel framework that distills the prior knowledge from a 2D Vision Transformer pre-trained on extensive 2D natural images to facilitate neuronal morphological learning of our 3D Vision Transformer. To bridge the knowledge gap between the 2D natural image and 3D microscopic morphologic domains, we propose a deformable tubular transferring strategy that adapts the pre-trained 2D natural knowledge to the inherent tubular characteristics of neuronal structure in the latent embedding space. The experimental results on the Janelia dataset of the BigNeuron project demonstrate that our method achieves a segmentation performance improvement of 4.53% in mean Dice and 3.56% in mean 95% Hausdorff distance.
@@ 1 Introduction
Reconstructing the 3D structure of single neurons plays a fundamental role in analyzing and understanding brain functionality and mechanisms [17, 6, 4, 7, 9, 11]. This intricate process traces the tree-like neuron morphology and digitizes the neuronal structure from light microscopic images. Nevertheless, this task has consistently remained challenging. Historically, neuron reconstruction heavily relied on manual annotation by neuroscience specialists, which requires significant time and effort [12]. Directly reconstructing neuronal morphology through various conventional tracing techniques often yields suboptimal outcomes due to the structural complexity of neurons, as well as the inherent noises and discontinuities present in light microscope imaging data.
@@
In this work, we design an innovative framework, named as DINeuro, to improve neuronal morphological learning by fully extracting the inherent tree-shape properties of neuron structures. This proposed method leverages the prior knowledge of DINO [1]—a 2D Vision Transformer pre-trained on vast 2D natural images—and transfers it to our 3D Vision Transformer to capture the topological and geometric features of 3D neuron images. To bridge the gap between 2D natural image and 3D neuronal topology, we design a deformable tubular transferring strategy to dynamically distill the 2D natural knowledge to the varied tubular characteristics of neuron morphology. In summary, our contributions are as follows:
@@ 2.2.2 Multi-View Tubular Adaptation and Iterative Tubular Offset
Neuronal structures are usually highly anisotropic with varying diameters and angles. To effectively extract diverse geometric and topological information from different perspectives of volumetric neuron images, as presented in Figure 2, we propose a multi-view learning strategy. This strategy flattens the transferred 2D convolutional kernel into three distinct 3D tubular kernels along the z-axis, y-axis, and x-axis respectively. Deformation offsets are then applied to shift the tubular kernels and align them with the tubular morphology of neurons, enabling better capture of neuronal branching structures. Taking the kernel transferred along the x-axis as an example, the transferred 2D convolutional kernel from DINO is inflated and flattened in the direction of the x-axis to form a 3D kernel in a tubular shape. The length of this tubular kernel is $16\times 16$, with the x-coordinates of the grids being continuous and fixed. The x-coordinates of the grids in this kernel can be represented as: $x_{0+h}$, where $x_{0}$ is the central grid and $h=\{-127,..,128\}$, denoting the horizontal distance of each grid from the central grid. While the y-coordinate and z-coordinate for each grid in this tubular convolution kernel are generated by iterative offsets. Through iteratively producing an offset based on the offset of the preceding grid point within the kernel and constraining the step of each offset, we ensure a coherent and continuous transformation of the kernel into a tubular shape and effectively minimize the risk of significant deviation. A grid position $(x_{i},y_{i},z_{i})$ is dependent on the previous grid $(x_{i-1},y_{i-1},z_{i-1})$ and can be described by: $y_{i}=y_{i-1}+\Delta y_{i},\quad z_{i}=z_{i-1}+\Delta z_{i}$, where $\Delta y_{i}$ and $\Delta z_{i}$ denote the offsets along the y-axis and z-axis for $y_{i}$ and $z_{i}$ respectively, ranging from $[-1,1]$. Therefore, the offsets for the y-coordinate and z-coordinate are iterative and accumulated, assuring the transferred 2D kernel is aligned with a continuous tubular neuron structure.
@@ 3.2 Results
To quantitatively evaluate the segmentation performance of our proposed framework, we employ two widely-used metrics in the field of image segmentation: mean Dice and mean 95% Hausdorff distance (Hd95). The SmartTracing method [2] is applied as the tracing algorithm to reconstruct neuron morphology from segmented results. Three spatial distance-driven metrics—Entire Structure Average (ESA), Different Structure Average (DSA), and Percentage of Different Structure (PDS)—are used to measure the geometric discrepancy between the reconstructed neuron and the manually annotated ground truth. The visualizations of segmentation results and reconstruction results for a sample are displayed in Figure 3.

### v2

@@ Abstract
Reconstructing neuron morphology from 3D light microscope imaging data is critical to aid neuroscientists in analyzing brain networks and neuroanatomy. With the boost from deep learning techniques, a variety of learning-based segmentation models have been developed to enhance the signal-to-noise ratio of raw neuron images as a pre-processing step in the reconstruction workflow. However, most existing models directly encode the latent representative features of volumetric neuron data but neglect their intrinsic morphological knowledge. To address this limitation, we design a novel framework, namely DINeuro, that distills the prior knowledge from DINO which pre-trained on extensive 2D natural images to facilitate neuronal morphological learning of our 3D Vision Transformer. To bridge the knowledge gap between the 2D natural image and 3D microscopic morphologic domains, we propose a deformable tubular transferring strategy that adapts the pre-trained 2D natural knowledge to the inherent tubular characteristics of neuronal structure in the latent embedding space. The experimental results on the Janelia dataset of the BigNeuron project demonstrate that our method achieves a segmentation performance improvement of 4.53% in mean Dice and 3.56% in mean 95% Hausdorff distance.
@@ 1 Introduction
Reconstructing the 3D structure of single neurons plays a fundamental role in analyzing and understanding brain functionality and mechanisms [9, 7, 11, 6, 4, 13]. This intricate process traces the tree-like neuron morphology and digitizes the neuronal structure from light microscopic images. Nevertheless, this task has consistently remained challenging. Historically, neuron reconstruction heavily relied on manual annotation by neuroscience specialists, which requires significant time and effort [14, 8]. Directly reconstructing neuronal morphology through various conventional tracing techniques often yields suboptimal outcomes [17] due to the structural complexity of neurons, as well as the inherent noises and discontinuities present in light microscope imaging data.
@@
In this work, we design an innovative framework that leverages the prior knowledge from DIN[1] to facilitate neuro morphological learning, namely DINeuro. This proposed method captures the inherent tree-shape properties of neuron structures. It is achieved by transferring the knowledge from 2D natural images into our 3D Vision Transformer to capture the topological and geometric features of 3D neuron images. To bridge the gap between 2D natural images and 3D neuronal topology, we design a deformable tubular transferring strategy to dynamically distill the 2D natural knowledge to the varied tubular characteristics of neuron morphology. In summary, our contributions are as follows:
@@ 2.2.2 Multi-View Tubular Adaptation and Iterative Tubular Offset
Neuronal structures are usually highly anisotropic with varying diameters and angles. To effectively extract diverse geometric and topological information from different perspectives of volumetric neuron images, as presented in Figure 2, we propose a multi-view learning strategy. This strategy flattens the transferred 2D convolutional kernel into three distinct 3D tubular kernels along the z-axis, y-axis, and x-axis respectively. Deformation offsets are then applied to shift the tubular kernels and align them with the tubular morphology of neurons, enabling better capture of neuronal branching structures. Taking the kernel transferred along the x-axis as an example, the transferred 2D convolutional kernel from DINO is inflated and flattened in the direction of the x-axis to form a 3D kernel in a tubular shape. The length of this tubular kernel is $16\times 16$, with the x-coordinates of the grids being continuous and fixed. The x-coordinates of the grids in this kernel can be represented as: $x_{0+h}$, where $x_{0}$ is the central grid and $h=\{-127,..,128\}$, denoting the horizontal distance of each grid from the central grid. The y-coordinate and z-coordinate for each grid in this tubular convolution kernel are generated by iterative offsets. Through iteratively producing an offset based on the offset of the preceding grid point within the kernel and constraining the step of each offset, we ensure a coherent and continuous transformation of the kernel into a tubular shape and effectively minimize the risk of significant deviation. A grid position $(x_{i},y_{i},z_{i})$ is dependent on the previous grid $(x_{i-1},y_{i-1},z_{i-1})$ and can be described by: $y_{i}=y_{i-1}+\Delta y_{i},\quad z_{i}=z_{i-1}+\Delta z_{i}$, where $\Delta y_{i}$ and $\Delta z_{i}$ denote the offsets along the y-axis and z-axis for $y_{i}$ and $z_{i}$ respectively, ranging from $[-1,1]$. Therefore, the offsets for the y-coordinate and z-coordinate are iterative and accumulated, assuring the transferred 2D kernel is aligned with a continuous tubular neuron structure.
@@ 3.2 Results
To quantitatively evaluate the segmentation performance of our proposed framework, we employ two widely used metrics in the field of image segmentation: mean Dice and mean 95% Hausdorff distance (Hd95). The SmartTracing method [2] is applied as the tracing algorithm to reconstruct neuron morphology from segmented results. Three spatial distance-driven metrics—Entire Structure Average (ESA), Different Structure Average (DSA), and Percentage of Different Structure (PDS)—are used to measure the geometric discrepancy between the reconstructed neuron and the manually annotated ground truth. The visualizations of segmentation results and reconstruction results for a sample are displayed in Figure 3.